OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3).

OpenStudy (anonymous):

1.substitute x=-1 in f(x)=x^2+1 ----> f(-1)=(-1)^2+1 2. substitute x= -3 in g(x)=3x+1 ----> g(-3)=3(-3)+1 3. in finding f(-1) + 5g(-3) just substitute what u had solved in 1 and 2

OpenStudy (anonymous):

f(-1)=(-1)^2+1=2 g(-3)=3(-3)+1=-9+1=-8 5g(-3)=5(-8)=-40 f(-1)+5g(-3)=2-40=-38